What is the slope point slope form of the equation of the line in slope equals 5 passing through 6 7
Answer by Boreal(15235)

y-y1=m(x-x1) is the point slope formula, where (x1,y1) are the point coordinates, and m the slope
y-7=5(x-6), Here, the signs are easier, but be careful when the point has a negative sign.
y-7=5x-30
y=5x-23
Check
7=5(6)-23, which works.